OpenStudy (anonymous):

A particle moves with the equation s(t)=t^2 -3t +2 where t is seconds and s(t) is meters A. When is the particle at rest B. what is the average velocity over the first 5 secs C. What is the average velocity on the interval [4,5] D. What is instantaneous velocity at 5 secs?

OpenStudy (anonymous):

the answer of a is 1.5 sekon.., V =0, V = ds(t)/dt = d/dt (t^2 -3t +2) = 2t -3; V = 2t -3. If V =0, so 0 = 2t-3; t=3/2 = 1.5 sekon please wait for other.,